Intercontinental, Inc., provides you with the following data for its single product:

Sales price per unit $ 52.00
Fixed costs (per month):
Selling, general, and administrative (SG&A) 1,350,000
Manufacturing overhead 2,700,000
Variable costs (per unit):
Direct labor 8.00
Direct materials 12.00
Manufacturing overhead 10.00
SG&A 6.00
Number of units produced per month 300,000 units


Required:

Compute the amounts for each of the following assuming that both production levels are within the relevant range. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

300,000 units 400,000 units
a. Prime cost per unit. $20.00 $20.00
b. Contribution margin per unit. $16.00 $16.00
c.

Gross margin per unit.

d. Conversion cost per unit.
e. Variable cost per unit.
f. Full absorption cost per unit.
g. Variable production cost per unit.
h. Full cost per unit.
Answer #1

Answer:

300000 Units 400000 Units
Amount($) Amount($)
1 Prime cost per unit 8+12                                     20.00 8+12                        20.00
2 Contribution margin per unit 52-8-12-10-6                                     16.00 52-8-12-10-6                        16.00
3 Gross margin per unit 52-8-12-10-9*                                     13.00 52-8-12-10-6.75**                        15.25
4 Conversion Cost per unit 8+10+9                                     27.00 8+10+6.75                        24.75
5 Variable cost per unit 8+12+10+6                                     36.00 8+12+10+6                        36.00
6 Full absorption Cost per unit 8+12+10+9                                     39.00 8+12+10+6.75                        36.75
7 Variable production cost per unit 8+12+10                                     30.00 8+12+10                        30.00
8 Full cost per unit 36+13.5***                                     49.50 36+10.125                      46.125
*9=27,00,000/300,000 units
**6.75=27,00,000/400,000 units
***13.50=27,00,000+13,50,000/300,000 units
****10.125=27,00,000+13,50,000/400,000 units